Dots sells T-shirt ($2) and shorts ($4). In April, total sales were $600. People bought 4 times as many T-shirt as shorts. How many T-shirt and shorts did Dots sell?
Answer by richwmiller(17219)

x+y=250,
2*x+4*y=600
x=250-y
2*(250-y)+4*y=600
500-2y+4*y=600
2*y=100
y=50
x=250-y
x=200 at 2 and y=50 at 4
check
200+50=250
2*x+4*y=600
2*200+4*50=600
400+200=600
600=600
ok